# GABAR: Graph Attention-Based Action Ranking for Relational Policy Learning

### v2

@@ Introduction
For action schema selection: the model needs to learn to assign the highest score to the correct action schema among all possible schemas. For object selection: for each parameter position, the model needs to learn to assign the highest score to the correct object among all candidate objects. This is done using the following loss function.
@@
Loss Function Given a training instance $(G,y_{a},y_{o})$, GABAR computes action scores $s_{a}$ for all possible action schemas and object scores $s_{o_{i}}$ for each parameter position $i$. The total loss is computed as $L=L_{action}+L_{objects}$, where, $L_{objects}$ is the sum of cross-entropy losses between $\text{softmax}(s_{o_{i}})$ and $y_{o_{i}}$ for each parameter position $i$.
@@
Training Procedure We train the model using the Adam optimizer with a learning rate of $0.0005$, $9$ rounds of GNN, and batch size of $16$. Training proceeds for a maximum of 500 epochs, and we select the model checkpoint that achieves the lowest loss on the validation set for evaluation.
Experiments
Table 1: Performance of both GABAR and GABAR-G methods across domains(number of problems shown in parenthesis(#)). Coverage indicates the percentage of test instances solved within 1000 steps. Plan lengths are reported only for solved instances.
GABAR GABAR-G
Domain Cov.(%) $\uparrow$ Mean PL $\downarrow$ Median PL $\downarrow$ PQR $\uparrow$ Cov.(%) $\uparrow$ Mean PL $\downarrow$ Median PL $\downarrow$ PQR $\uparrow$
Blocks World (100) 100 53.1 46 1.42 84 52.6 54.7 1.1
Gripper (100) 100 147.1 133 0.99 89 146.7 137 0.95
Miconic (50) 100 181.7 181 0.97 75 192.4 189 0.87
Logistics (50) 76 155.2 132 0.65 12 238.5 220 0.33
Visitall (50) 88 207.9 190 1.07 63 205.7 183 0.86
Grid (50) 96 72.1 62 0.81 68 70.2 68 0.73
@@
We evaluate GABAR’s performance across a diverse set of classical planning domains. Our experiments aim to assess both the quality of learned policies and their ability to generalize to significantly larger problems than those in training.
Domains. We selected six standard planning domains that present different types of structural complexity and scaling dimensions.
@@
Logistics involves transporting packages between locations using trucks (for intra-city transport) and airplanes (for inter-city transport). The domain scales with both the number of cities (4-8 for training/validation, 15-30 for testing) and packages (3-10 for training/validation, 9-24 for testing).
@@
Grid involves navigating through a grid where certain doors are locked and require specific keys to open. The number of locks (3) and keys (5) remain the same across training and testing while varying the size of the grid (7$\times$9 for training/validation to 11$\times$14 for testing - a 150% increase in cells).
Evaluation Metrics. We evaluate GABAR using three primary metrics:
  • Coverage (Cov): The percentage of test instances successfully solved within a 1000-step limit.
  • Plan Length (PL): Measured through both mean and median plan lengths across solved instances.
  • Plan Quality Ratio(PQR): Ratio of plan length produced by Fast Downward (FD) planner to the plan length produced by the learned policy. FD is run with fd-lama-first setting. We chose this satisficing configuration over optimal planners since optimal planners fail to solve most test problems within a reasonable time. While this means we cannot guarantee the optimality of the reference plans, it provides a practical baseline for assessing solution quality across our test suite.
High coverage on larger instances demonstrates the model’s ability to learn robust action selection strategies, while plan length and quality metrics reveal whether these strategies remain efficient as problems scale up.
Ablation: To understand the importance of different architectural components, we conduct an ablation study focusing on the design choice of using global nodes in the graph (GABAR-G). This variant of GABAR is without the global node, which helps assess its role in information propagation and helps quantify how the global node affects coverage, plan quality, and length.
Results and Discussions
Table 1 shows the results of our learning system across $6$ domains. We can see that the proposed method (OURS) generalizes very well on three domains - Blocksworld, Gripper, and Miconic. We can also see that our method generalizes well in the other three domains. This result, while not as good as others, shows an important capability of the proposed system - the ability to handle complex domains.
Both Grid and Logistics are more complex domains (Ståhlberg, Bonet, and Geffner 2022b), as expressing properties of the environment requires multiple relations (Ex, the feature expressing that a package is in a city while possibly within a vehicle needs multiple relations to express). We can see that the proposed method can learn these complex features that are required for solving the planning problem as well as generalize to larger problems in these settings (success rate of 76% and 96% in them). While the success rate of Visitall is not as impressive as Blocksworld, the generalization it achieves is impressive - it is solving problems up to eight times larger than what it trained on with an 88% success rate.
The ablation study comparing GABAR with and without the global node (GABAR vs. GABAR-G) reveals that removing the global node leads to degradation in performance.
In domains that primarily require local reasoning, like Blocks World, the impact is moderate but still notable - coverage drops from 100% to 84%, and the plan quality ratio decreases from 1.42 to 1.1. This suggests that while local messages passing between objects and actions can capture basic patterns, the global node helps maintain a broader context. This degradation is even more dramatic in Logistics, where coverage plummets from 76% to just 12%. The stark difference can be attributed to the need for long-range planning, as logistics requires orchestrating multiple vehicles across different cities.
Plan quality also suffers significantly without the global node. In Logistics, the plan quality ratio drops from 0.65 to 0.33, indicating that solutions become nearly twice as inefficient. This degradation in quality suggests that the global node plays a crucial role in helping the model learn strategic action selection rather than just locally reasonable choices.
These findings validate our architectural choice of including a global node in GABAR. The global node proves essential for facilitating long-range information sharing that helps coordinate actions across distant entities.
Conclusion and Future Work
We presented GABAR, a novel graph-based architecture for learning generalized policies in classical planning through action ranking. Our key contributions include (1) an action-centric graph representation that explicitly captures action-object relationships, (2) a GNN architecture augmented with global nodes and GRUs for effective information propagation, and (3) a sequential decoder that learns to construct complete grounded actions.
Our experimental results demonstrate strong generalization capabilities across multiple planning domains. GABAR achieves great coverage on Blocksworld, Gripper, Visitall, and Miconic domains when scaling to problems that are many times larger than those used in training. The architecture also shows promising results in more complex domains like Grid and Logistics. The ablation study removing global nodes highlights their critical role in enabling effective information propagation.
The primary scaling challenge lies in the graph representation’s growth rate, particularly for domains with high-arity actions or predicates. Future work could explore more compact representations while maintaining expressiveness and investigate ways for pruning irrelevant actions in the graph.
Acknowledgments

### v3

@@ 3.4 Data Generation and Training
For action schema selection, the model needs to learn to assign the highest score to the correct action schema among all possible schemas. For each parameter position, the model needs to learn to assign the highest score to the correct object among all candidate objects. This is done using the following loss function.
@@
Loss Function. Given a training instance $(G,y_{a},y_{o})$, GABAR computes action scores $s_{a}$ for all possible action schemas and object scores $s_{o_{i}}$ for each parameter position $i$. The total loss is computed as $L=L_{action}+L_{objects}$, where, $L_{objects}$ is the sum of cross-entropy losses between $\text{softmax}(s_{o_{i}})$ and $y_{o_{i}}$ for each parameter position $i$.
@@
Training Procedure. For all domains, we train the model using the Adam optimizer with a learning rate of $0.0005$, 99 rounds of GNN, and batch size of $16$, hidden dimensionality of $64$. Training proceeds for a maximum of 500 epochs, and we select the model checkpoint that achieves the lowest loss on the validation set for evaluation. It takes between 1-2 hours to train a model for each domain on an RTX 3080.
4 Experiments
@@
We evaluate GABAR’s performance across a diverse set of classical planning domains. Our experiments aim to assess both the quality of learned policies and their ability to generalize to significantly larger problems than those in training. We selected eight standard planning domains that present different types of structural complexity and scaling dimensions. The domains are : BlocksWorld, Gripper, Miconic, Spanner, Logistics, Rovers, Grid and Visitall (More details in appendix A.1).
We divide the test set for each domain into 3 separate subsets, easy, medium, and hard with increasing difficulty with problem sizes as defined in table 1 along with the training and validation dataset sizes. Each test subset has 100 problems. In contrast, the training dataset consists of problems smaller and simpler than the ones found in the easy subset.
Table 1: Distribution of problem sizes across train, validation, and test datasets. The ranges indicate the number of objects/variables defining each domain’s problem complexity.* In Visitall and Grid, size refers to number of cells
domain Train Val Test domain Train Val Test
easy med hard easy med hard
Blocks $[6,9]$ $[10]$ $[11,20]$ $[21,30]$ $[31,40]$ Rovers $[3,9]$ $[10]$ $[11,30]$ $[31,50]$ $[51,70]$
Gripper $[5,15]$ $[17]$ $[20,40]$ $[41,60]$ $[61,100]$ Visitall* $[9,36]$ $[49]$ $[50,100]$ $[101,200]$ $[201,400]$
Miconic $[1,9]$ $[10]$ $[20,40]$ $[41,60]$ $[61,100]$ Grid* $[25,49]$ $[63]$ $[64,100]$ $[100,125]$ $[125,154]$
Spanner $[2,9]$ $[10]$ $[11,20]$ $[21,30]$ $[31,40]$ Logistics $[4,7]$ $[8]$ $[15,20]$ $[21,25]$ $[26,30]$
We evaluate GABAR against three state-of-the-art approaches for learning generalized policies: GPL, ASNets, and GRAPL. Each represents a different architectural paradigm for tackling generalized planning.
4.1 Baseline Approaches
GPL (Generalized Policy Learning) unsupervised_gnn learns value functions over states using GNNs. It selects actions by identifying unvisited states with the lowest estimated cost-to-goal. GPL attempts to learn a globally consistent value function by minimizing the regularized Bellman error $|V(s)-(1+\min_{s^{\prime}\in N(s)}V(s^{\prime}))|+\max\{0,V^{*}(s)-V(s)\}+\max\{0,V(s)-\delta V^{*}(s)\}$ where $s$ is state and $V(s)$ is value function defined as cost to goal.
ASNets (Action Schema Networks) toyer2020asnets employs a neural network with alternating action and proposition layers. Each layer contains modules that connect only to related modules in adjacent layers. ASNets shares weights across modules of the same action schema or predicate, enabling generalization across problems of varying sizes but limiting its reasoning to a fixed-depth receptive field.
GRAPL (Graph Relational Action Ranking for Policy Learning) karia2021learning learns to rank actions using canonical abstractions where objects with identical properties are grouped together. Its neural network predicts both action probabilities and plan length estimates but lacks the explicit action-object graph structure and conditional decoding mechanisms of our approach.
Ablations: To assess each component’s contribution to GABAR, we conducted ablation experiments with three variants:
  • GABAR-ACT: Removes action nodes and action-object edges from the graph, using only object and predicate nodes for decision-making.
  • GABAR-CD: Eliminates conditional decoding by selecting action parameters independently rather than sequentially. This tests the importance of dependencies between action parameters.
  • GABAR-RANK: Replaces action ranking with a value function learning objective while maintaining our graph representation.
These ablations isolate our key innovations—action-centric graphs, GRU-based conditional decoding, and action ranking—to measure their individual impact on performance.
Evaluation Metrics. We evaluate GABAR using two primary metrics:
  • Coverage (C): The percentage of test instances successfully solved within a 1000-step limit.
  • Plan Quality Ratio(P): Ratio of plan length produced by Fast Downward (FD) (fast_downward, ) planner to the plan length produced by the learned policy. FD is run with fd-lama-first setting. We chose this satisficing configuration over optimal planners since optimal planners fail to solve most test problems within a reasonable time. While this means we cannot guarantee the optimality of the reference plans, it provides a practical baseline for assessing solution quality.
High coverage on larger instances demonstrates the model’s ability to learn general action selection strategies, while plan quality ratio reveals whether these strategies produce efficient plans as problems scale up.
5 Results and Discussions
Table 2: Coverage (C) and plan quality ratio (P) across different domains and difficulty levels. Turqoise color intensity for C values indicates coverage score (Only values above 50% are highlighted: Light: 50-74%, Medium: 75-89%, Dark: 90-100%). P values are highlighted in violet with similar thresholds (No highlight: 0-0.49, Light: 0.5-0.74, Medium: 0.75-0.99, High: 1.0+). Columns are grouped into Baselines and Ablations. Combined rows show averages across all domains. (*) only non-zero values from their respective domains were considered
Baselines Ablations
GPL ASNets GRAPL GABAR GABAR-ACT GABAR-CD GABAR-RANK
[…]
Blocks E 100 1.1 100 1.6 64 0.65 100 1.5 44 0.65 100 0.92 29 0.79
M 45 0.68 100 1.5 48 0.44 100 1.6 14 0.49 92 0.81 21 0.71
H 10 0.33 92 1.4 38 0.28 100 1.7 4 0.35 81 0.80 9 0.61
Miconic E 97 0.97 100 1.0 68 0.56 100 1.0 35 0.55 97 0.88 42 0.67
M 37 0.56 100 0.98 65 0.54 100 0.97 18 0.33 94 0.86 29 0.37
H 19 0.29 90 0.92 60 0.49 100 0.95 2 0.27 88 0.83 16 0.29
Spanner E 73 1.1 78 0.86 22 0.65 94 1.1 31 0.65 87 0.98 57 0.82
M 42 0.56 60 0.69 5 0.55 93 0.99 11 0.27 81 0.93 42 0.77
H 3 0.18 42 0.61 0 - 89 0.91 0 - 62 0.79 12 0.45
Gripper E 100 1.0 78 0.98 26 0.95 100 1.1 31 0.56 95 1.0 55 0.58
M 56 0.85 54 0.91 12 0.67 100 0.99 23 0.40 92 0.93 43 0.41
H 21 0.74 42 0.88 0 - 100 0.96 9 0.28 87 0.86 21 0.33
Visitall E 69 1.3 94 0.96 92 1.1 93 1.1 72 1.2 91 1.1 52 0.64
M 15 0.76 86 0.93 88 1.0 91 1.0 64 0.93 89 1.1 46 0.56
H 0 0 64 0.81 78 0.99 88 1.1 44 0.67 83 1.2 39 0.54
Grid E 74 0.89 52 0.81 20 0.38 100 0.91 21 0.56 79 0.87 17 0.54
M 17 0.61 45 0.66 3 0.28 97 0.85 8 0.46 71 0.65 12 0.28
H 0 0 21 0.60 0 - 92 0.74 0 - 54 0.53 0 -
Logistics E 56 0.61 39 0.71 32 0.81 90 0.75 12 0.64 31 0.86 41 0.65
M 7 0.21 22 0.55 9 0.45 76 0.65 3 0.49 25 0.54 21 0.49
H 0 0 4 0.39 0 - 71 0.59 0 - 6 0.35 0 -
Rovers E 64 0.99 67 0.96 21 0.35 87 1.0 22 0.75 44 0.81 33 0.67
M 9 0.32 56 0.87 5 0.19 82 0.96 6 0.66 37 0.63 9 0.56
H 0 0 31 0.64 0 - 77 0.97 0 - 19 0.57 0 -
Combined E 79.1 0.98 76 0.98 43.5 0.67 95.5 1.0 33.5 0.69 78 0.93 40.2 0.67
M 28.5 0.56 65.4 0.88 29.3 0.51* 92.2 1.0 18.4 0.50 72.7 0.80 27.8 0.51
H 6.5 0.39* 48.5 0.78 22.1 0.58* 89.2 0.99 7.4 0.39* 60 0.73 12.1 0.44*
Table 2 shows the performance of GABAR, its ablations, and the baseline approaches across all domains. The results demonstrate significant differences in generalization capabilities and solution quality.
5.1 Generalization Performance
GABAR generalizes remarkably well across all domains, with coverage dropping minimally from 95.5% on easy problems to 89.5% on hard problems. This generalization is particularly evident in domains like Visitall, where GABAR solves problems 8 times larger than those in training. On BlocksWorld, Gripper, and Miconic, GABAR achieves 100% success rate at all difficulty levels.
Even in complex domains requiring multi-relation features (Grid, Logistics) or higher arity predicates (Rovers), GABAR maintains strong performance with success rates of 92%, 62%, and 77% respectively on hard instances. This demonstrates that our graph structure effectively encodes critical problem information, and the decoder successfully captures relationships between actions and their parameters.
We now analyze GABAR against baselines and relevant ablation results, highlighting how our design choices contribute to performance improvements.
Action Ranking vs. Value Function Learning: GPL uses GNNs to learn value functions over states, aiming to produce globally consistent estimates of cost-to-goal unsupervised_gnn . As problems grow larger, maintaining these consistent estimates becomes significantly more challenging. GPL’s coverage drops to 0% on hard instances of Logistics and Visitall where GABAR maintains 71% and 88% coverage. Our ablation GABAR-RANK, which replaces action ranking with value function learning while keeping our graph structure intact, shows a similar pattern—coverage decreases from 89.2% to just 12.1% on hard problems. This parallel between GPL and GABAR-RANK confirms that directly ranking actions is more effective than learning value functions.
Conditional vs. Non-conditional Decoding: GRAPL also learns action rankings but uses an abstract state representation without explicit modeling of action parameter dependencies karia2021learning . While conceptually similar to our approach, GRAPL achieves only moderate performance on easy instances (43.5% coverage) and struggles on harder problems (22.1% coverage) compared to GABAR’s 95.5% and 89.2%. Our ablation GABAR-CD, which removes our conditional decoding mechanism, shows a similar degradation—coverage drops from 89.2% to 60.0% on hard problems overall, with particularly dramatic decreases in complex domains like Logistics (from 62% to 21%). This comparison highlights the critical importance of our conditional GRU-based decoder for handling complex dependencies in domains where objects participate in multiple predicates simultaneously (e.g., a package being in both a city and a vehicle). The decoder learns relationships between parameter selections by conditioning each object choice on previously selected objects, enabling the model to capture complex inter-parameter relationships that GRAPL cannot.
Action-Centric Graph Structure: ASNets employs alternating action and proposition layers with a fixed-depth receptive field toyer2020asnets , limiting its ability to reason about long chains of dependencies. While ASNets performs reasonably on medium-difficulty problems, it generalizes poorly to larger instances—achieving only 21%, 4%, and 31% coverage on hard instances of Grid, Logistics, and Rovers, compared to GABAR’s 92%, 62%, and 77%. Although we don’t have a perfect ablation parallel for ASNets, our GABAR-ACT ablation (removing action nodes) shows catastrophic performance drops from 95.5% to 33.5% on easy problems and from 89.2% to just 7.4% on hard problems. This suggests that explicit action-object relationship modeling in the graph structure is critical for robust generalization.
These comparisons demonstrate that GABAR’s superior performance stems from the synergistic combination of three key ideas: (1) an action-centric graph representation that captures relationships between actions and objects, (2) conditional decoding that models dependencies between action parameters, and (3) direct action ranking that avoids the challenges of learning globally consistent value functions. The ablation studies confirm that each component addresses these limitations present in current approaches to generalized planning.
6 Conclusion and Future Work
We presented GABAR, a novel graph-based architecture for learning generalized policies in classical planning through action ranking. Our key contributions include (1) an action-centric graph representation that explicitly captures action-object relationships, (2) a GNN architecture augmented with global nodes and GRUs for effective information propagation, and (3) a sequential decoder that learns to construct complete grounded actions. First, GABAR shows strong generalization capabilities, successfully solving problems significantly larger than those used in training. The system maintains high coverage (89%) even on the hardest test instances that are up to 8 times larger than training problems. This generalization is particularly significant in complex domains like logistics and grid that require coordinating multiple entities. Second, our comparison with GPL demonstrates that learning to rank actions is more generalizable and sample-efficient than learning value functions. GABAR achieves superior performance using only 25% of the training data required by GPL while maintaining better coverage and plan quality across all difficulty levels. Third, our ablation study validates the importance of the global node in facilitating long-range information sharing. Without this architectural component, performance degrades significantly on larger problems, particularly in domains requiring complex coordination. The primary challenge for the proposed method is the graph representation’s growth rate, particularly for domains with high-arity actions or predicates. Future work could explore more compact representations while maintaining expressiveness and investigate ways of pruning irrelevant actions in the graph. Other future work could also include expanding the representation to handle uncertainty and solve more real-world problems such as object rearrangement.
7 Acknowledgments
Appendix A Appendix
A.1 Domain descriptions
@@
Logistics involves transporting packages between locations using trucks (for intra-city transport) and airplanes (for inter-city transport). The domain scales with both the number of cities (4-8 for training/validation, 15-30 for testing) and packages (3-9 for training/validation, 10-24 for testing).
@@
[…]
Spanner requires tightening nuts at one end of a corridor with spanners collected along the way. Number of locations varies from 10-20 in training and testing, the number of spanners varies from 2-10 in training and 11 to 40 in testing.
Rovers requires multiple rovers equipped with different capabilities to perform experiments and send results back to the lander. Training and validation instances use 2-5 rovers and 3-10 waypoints; those for testing have 3-7 rovers and 11-70 waypoints.
Table 3 contains information about number of datapoints used for each domain  [^1] Code : https://anonymous.4open.science/r/ltp-4741/ . Dataset submitted as supplementary material. All problems were generated using openly available PDDL-generators [17].
Table 3: Number of training datapoints per domain
Domain Training Datapoints
Blocksworld 3,348
Gripper 6,287
Miconic 4,458
Spanner 3,801
Logistics 6,556
Rovers 3,874
VisitAll 3,654
Grid 6,880
A.2 Ablations
We present more ablations on our method in this section to further analyze the contribution of different components in our GABAR architecture. The two configurations tested are:
  • GABAR-G: An ablated version of GABAR that removes the global node from the graph neural network. Without the global node, the model lacks a centralized representation that aggregates information across the entire state, potentially limiting its ability to reason about complex relationships spanning multiple objects.
  • GABAR-ACT_CD: A severely limited version that removes both the action representation from the graph and the conditional decoding mechanism. This baseline lacks the capability to explicitly represent actions in the state graph and cannot condition parameter selection on previously chosen actions, effectively testing the importance of both structured action representation and sequential decision-making. Results Discussion
The results from Table 4 clearly demonstrate the effectiveness of our complete GABAR architecture compared to its ablated counterparts:
Overall Performance: GABAR significantly outperforms both ablated versions across all domains and difficulty levels. With average coverage of 95.5%, 92.2%, and 89.2% for easy, medium, and hard problems respectively, GABAR demonstrates robust generalization capabilities.
GABAR v/s GABAR-G : The dramatic performance drop from GABAR to GABAR-G (especially on hard problems where coverage falls from 89.2% to 42.5%) highlights the critical importance of the global node in our architecture. This component enables effective information aggregation across the entire state graph, substantially improving the model’s ability to handle complex planning scenarios. This degradation in quality suggests that the global node plays a crucial role in helping the model learn strategic action selection rather than just locally reasonable choices.
GABAR v/s GABAR-ACT_CD : The extremely poor performance of GABAR-ACT_CD (with coverage below 10% in most domains and near-zero on harder problems) confirms that both action representation within the graph and conditional decoding are essential components. Without them, the model fails to capture the relationship between actions and their parameters or to maintain consistency in sequential decisions.
Plan Quality: Beyond just solving more problems, GABAR consistently generates higher-quality plans (as measured by PQR) compared to its ablated counterparts. This indicates that the full model not only finds solutions but discovers more efficient paths to the goal.
Table 4: Performance across domains and difficulty levels. Coverage (C) indicates percentage of test instances solved within 1000 steps. Plan Quality Ratio (P) reported only for solved instances. Turqoise color intensity for C values indicates coverage score (Only values above 50% are highlighted: Light: 50-74%, Medium: 75-89%, Dark: 90-100%). P values are highlighted in violet with similar thresholds (No highlight: 0-0.49, Light: 0.5-0.74, Medium: 0.75-0.99, High: 1.0+). Combined rows show averages across all domains. (*) only non-zero values from respective domains considered
ACT_CD GABAR GABAR-G
[…]
Blocks E 11 0.33 100 1.5 100 1.2
M 8 0.31 100 1.6 88 1.2
H 0 - 100 1.7 32 0.70
Miconic E 8 0.32 100 1.0 95 0.95
M 3 0.27 100 0.97 70 0.87
H 0 - 100 0.95 56 0.73
Spanner E 12 0.43 94 1.1 82 0.93
M 1 0.24 93 0.99 74 0.81
H 0 - 89 0.91 54 0.62
Gripper E 15 0.58 100 1.1 95 1.0
M 5 0.34 100 0.99 92 0.95
H 0 - 100 0.96 77 0.90
Visitall E 23 0.58 93 1.1 77 0.96
M 19 0.55 91 1.0 66 0.83
H 16 0.48 88 1.1 29 0.51
Grid E 0 - 100 0.91 79 0.83
M 0 - 97 0.85 68 0.62
H 0 - 92 0.74 33 0.41
Logistics E 0 - 90 0.75 45 0.43
M 0 - 76 0.65 22 0.33
H 0 - 71 0.59 4 0.21
Rovers E 3 0.45 87 1.0 77 0.85
M 0 - 82 0.96 72 0.79
H 0 - 77 0.97 55 0.65
Combined E 9 0.44* 95.5 1.0 80.2 0.89
M 4.5 0.34* 92.2 1.0 69.2 0.79
H 2 0.48* 89.2 0.99 42.5 0.59
A.3 Beam Search Action Decoder
The beam search action decoder (Algorithm 2) enhances the basic greedy decoder by maintaining multiple candidate sequences at each decoding step.
Initialization (line 1): Similar to the greedy approach, the decoder initializes its hidden state using the global graph embedding and a zero vector.
Action Schema Selection (lines 2-5): Instead of selecting only the highest-scoring action schema, beam search maintains the top-$k$ action schemas. For each action schema $a\in A$, a score $s_{a}$ is computed (line 3). The $k$ highest-scoring actions form the initial beam $B_{0}$ (lines 4-5), each representing a partial sequence with an action schema, an empty parameter list, the initial hidden state, and the action score.
Parameter Selection with Beam Search (lines 6-15): During each decoding step, the algorithm expands all sequences in the current beam by considering the top-$k$ object candidates for the next parameter position:
  • For each sequence in the current beam $B_{t-1}$ (line 9), scores are computed for all objects using the current hidden state (line 10).
  • The top-$k$ objects $O^{*}$ are identified (line 11).
  • Each sequence is expanded with each of these top-$k$ objects (lines 12-14), creating $k\times|B_{t-1}|$ new candidate sequences in $B^{\prime}_{t}$.
  • The hidden state is updated for each new sequence (line 13).
  • The beam is pruned to keep only the $k$ highest-scoring expanded sequences (line 15), forming the new beam $B_{t}$.
Output (line 16): After processing all parameter positions, the algorithm returns the set of completed sequences, each forming a fully grounded action.
In our implementation, we set $k=2$ for the beam search. This parameter choice provides a reasonable trade-off between exploration and computational efficiency. A beam width of 2 allows the decoder to maintain alternative sequences when faced with uncertainty, which is particularly important for handling invalid actions that might need to be discarded. Setting a larger beam width provides better results theoretically but increases computational cost without proportional benefits in our specific domain.
[…]
[0.8ex]
1: $h_{0}=\text{GRU}(\mathbf{g}^{L},\mathbf{0})$
[0.8ex] 2: // Step 1: Action schema selection
[0.5ex] 3: $s_{a}=\text{MLP}(h_{0}\odot\mathbf{v}^{L}_{a}),\forall a\in A$
[0.5ex] 4: $A^{*}=\{a_{1},a_{2},\ldots,a_{k}\}=\underset{a\in A}{\text{top-}k}(s_{a})$
[0.5ex] 5: $B_{0}=\{(a_{j},[],h_{0},s_{a_{j}})|a_{j}\in A^{*}\}$
[0.8ex] 6: // Step 2: Parameter selection with beam search
[0.5ex] 7: for $t=1$ to max_params do
[0.5ex] 8:   $B^{\prime}_{t}=\emptyset$
[0.5ex] 9:   for $(a,[o_{1},\ldots,o_{t-1}],h_{t-1},s)\in B_{t-1}$ do
[0.5ex] 10:    $s_{o}=\text{MLP}(h_{t-1}\odot\mathbf{v}^{L}_{o}),\forall o\in O$
[0.5ex] 11:    $O^{*}=\{o_{1},o_{2},\ldots,o_{k}\}=\underset{o\in O}{\text{top-}k}(s_{o})$
[0.5ex] 12:    for $o_{j}\in O^{*}$ do
[0.5ex] 13:    $h_{t}=\text{GRU}(h_{t-1},\mathbf{v}^{L}_{o_{j}})$
[0.5ex] 14:    $B^{\prime}_{t}=B^{\prime}_{t}\cup\{(a,[o_{1},\ldots,o_{t-1},o_{j}],h_{t},s+s_{o_{j}})\}$
[0.5ex] 15:   $B_{t}=\underset{b\in B^{\prime}_{t}}{\text{top-}k}(\text{score}(b))$
[0.8ex] 16: return $\{(a,[o_{1},o_{2},\ldots,o_{n}])|(a,[o_{1},o_{2},\ldots,o_{n}],h,s)\in B_{\text{max\_params}}\}$
Algorithm 2 Beam Search Action Decoder